OpenStudy (anonymous):

Determine the Descriminate of each equation, how many solutions does each equation have: A: 4x^2-2x=10 B: 3x^2+3=6x

hero (hero):

Hint: 1. If the discriminant is negative, then you have no solutions 2. If the discriminant is zero, then you have one solution. 3. If the discriminant is positive, then you have two solutions

hero (hero):

Make sure you put the quadratic equations in standard form first before calculating the discriminant.

OpenStudy (anonymous):

okay so 4x^2-2x -10=0

hero (hero):

Yes, so are you going to try calculating the discriminant now?

OpenStudy (anonymous):

so 2^2 - 4(4)(-10) 4 - 16(-10) 4 + 160 164?

hero (hero):

Be careful. The b term is negative, not positive. (-2)^2 - 4(4)(-10) 4 - 16(-10) 4 + 160 164

OpenStudy (anonymous):

oh okay, I should have paid more attention to that. so the descriminant is 164 and their are two solutions

hero (hero):

Also, you should probably get used to writing your steps in this manner: (-2)^2 - 4(4)(-10) = 4 - 16(-10) = 4 -(-160) <---- This step shows the proper property of a positive number = 4 + 160 = 164

OpenStudy (anonymous):

okay will do, so the next one is 3x^2 + 3 = 6x 3x^2 - 6x + 3 = 0 (-6)^2 + 4(3)(3) 36 + 12(3) 36 + 36 72, two solutions right?

hero (hero):

Looks right

OpenStudy (anonymous):

okay thank you